Find the derivative of the function. f(x) = arccsc 8x -1 f(x) = x\V 16x2 – 1 10 nyFind an equation of the tangent line to the graph of the function at the given point. y = - arccos y 37 2 xt. = - arccOS X
Answer #1

① f(거) Cosee (81) we know that I covecx : 1 xvxezi So f(x) = -1 - 8 81 faya | -1 Vu가지 ( ② 9 = as (x) ” 1 | VIX ( ) mNow equ of line passing through (x, ,9.) with n tangent. y-y, = m(21-2012 hese ms -,4,5 y, a so you can 11-6-) y (**) y egy =

I hope this helps.

thumbs up...plz

thanks